# Generalized Petermann factor of non-Hermitian systems at exceptional points

### v1

@@ Appendix C Fitting procedure
The goal in this Appendix is to determine all parameters in the effective Hamiltonian (100) from the numerically computable quantities $\Omega$ and $\psi$ that solve the mode equation (111). This is a non-trivial task as all parameters in $H$ are complex such that their phase cannot be fixed by comparing the intensity pattern of the eigenstates in the cavities. However, the phase of the parameters enter in the calculation for the response strength and therefore needs to be determined correctly. To do so, we employ a two-step fitting procedure. First, proper basis wavefunctions are constructed to project the modes $\psi_{1-4}$ into a $\mathbb{C}^{4}$ vector space. Second, a direct comparison of the projected modes to the eigenstates of the effective Hamiltonian (100) allows for a unique fit of the parameters.
@@ C.1 Auxiliary systems for basis modes
To do so, we design artificial systems which support pure CW and CCW propagating waves in each cavity. This is achieved by placing a single cavity–either cavity a or b–next to a waveguide. A single mirror of width $d=0.06R$ in the waveguide is then used to tune the system to an exceptional point via mirror-induced asymmetric backscattering [55, 56]. Thus, the mode is either a pure CCW or CW propagating wave depending on the mirror being left or right to the cavity. This mode is used as the respective basis state $\phi_{i}$ as shown in Fig. 4.
@@
With the fixed set of basis states $\phi_{1-4}$ each modes of the system $\psi_{j}$ can be mapped to a vector $a^{(j)}\in\mathbb{C}^{4}$ with components calculated from the overlap integral
@@
where the integration area $M$ is the area of the cavity where the basis mode is constructed from. Each mode $\psi_{j}$, in principle, can be multiplied by an arbitrary phase. To remove this ambiguity, the phase of each vector $a^{(j)}$ is specified to fulfill
@@ C.2 Calculation of the effective Hamiltonian
is used that compares the eigenstates $v^{(j)}$ of the Hamiltonian to the previously calculated representation of the mode $a^{(j)}$ as complex vectors. The minimizing of the cost function is done with a standard algorithm. The obtained values $t=10^{-4}(4.674614-2.915369i)$ and $r_{\text{b}}=10^{-4}(-1.541910-1.943754i)$ are unique in the sense that different initial guesses for the cost function optimization lead to the same parameter values.

### v2

@@ Appendix C Fitting procedure
The goal in this Appendix is to determine all parameters in the effective Hamiltonian (104) from the numerically computable quantities $\Omega$ and $\psi$ that solve the mode equation (115). This is a non-trivial task as all parameters in $H$ are complex such that their phase cannot be fixed by comparing the intensity pattern of the eigenstates in the cavities. However, the phase of the parameters enter into the calculation for the response strength and therefore needs to be determined correctly. To do so, we employ a two-step fitting procedure. First, proper basis wavefunctions are constructed to project the modes $\psi_{1-4}$ into a $\mathbb{C}^{4}$ vector space. Second, a direct comparison of the projected modes to the eigenstates of the effective Hamiltonian (104) allows for a unique fit of the parameters.
@@ C.1 Auxiliary systems for basis modes
To do so, we design artificial systems that support pure CW and CCW propagating waves in each cavity. This is achieved by placing a single cavity–either cavity a or b–next to a waveguide. A single mirror of width $d=0.06R$ in the waveguide is then used to tune the system to an exceptional point via mirror-induced asymmetric backscattering [59, 54]. Thus, the mode is either a pure CCW or CW propagating wave depending on the mirror being left or right to the cavity. This mode is used as the respective basis state $\phi_{i}$ as shown in Fig. 4.
@@
With the fixed set of basis states $\phi_{1-4}$ each mode of the system $\psi_{j}$ can be mapped to a vector $a^{(j)}\in\mathbb{C}^{4}$ with components calculated from the overlap integral
@@
where the integration area $M$ is the area of the cavity from which the basis mode is constructed. Each mode $\psi_{j}$, in principle, can be multiplied by an arbitrary phase. To remove this ambiguity, the phase of each vector $a^{(j)}$ is specified to fulfill
@@ C.2 Calculation of the effective Hamiltonian
is used that compares the eigenstates $v^{(j)}$ of the Hamiltonian to the previously calculated representation of the mode $a^{(j)}$ as complex vectors. The minimization of the cost function is done with a standard algorithm. The obtained values $t=10^{-4}(4.674614-2.915369i)$ and $r_{\text{b}}=10^{-4}(-1.541910-1.943754i)$ are unique in the sense that different initial guesses for the cost function optimization lead to the same parameter values.